y''-4y'+4y=(x+1)e^x


1
Expert's answer
2021-12-16T17:16:25-0500

y" - 4y'+4y = (x+1)ex

For complementary function , the auxiliary equation is

m² - 4m + 4 = 0

=> m² - 2.m.2 + 2² = 0

=> (m-2)² = 0

=> m = 2, 2

So the complementary function is given by

C.F. = (c1+c2x)e2x

Now let's find particular integral

P.I. = "\frac{1}{f(D)}(x+1)e^{x}" , where f(D) = (D-2)², Dn = "\frac{d^{n}}{dx^{n}}"

So P.I. = ex"\frac{1}{f(D+1)}(x+1)"

= ex "\frac{1}{(D+1-2)²}(x+1)"

= ex"\frac{1}{(D-1)²}(x+1)"

= ex"\frac{1}{(1-D)²}(x+1)"

= ex (1- D)-2(x+1)

= ex (1 + 2D + 3D² + •••••• upto ∞)(x+1)

= ex{(x+1) + 2(1) + 3(0) ••••• upto ∞}

= ex(x+3)

= (x+3)ex

So the general solution is y = C.F. + P.I.

That means general solution is

y = (c1+c2x)e2x + (x+3)ex where c1, and c2 are arbitrary constants.